OpenStudy (courtney_celiene1236):

1. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) =

OpenStudy (courtney_celiene1236):

@mayankdevnani

OpenStudy (sshayer):

\[t _{n}=4n \left( 4n+2 \right)=16n^2+8n\] \[\Sigma~t _{n}=16 ~\Sigma n^2+8 ~\Sigma n\]

OpenStudy (sshayer):

do you know \[\Sigma n^2 ~and~\Sigma ~n\]?

OpenStudy (sshayer):

\[\Sigma n^2=?\]

OpenStudy (courtney_celiene1236):

No. The lesson for this confused me entirely.

OpenStudy (sshayer):

\[\Sigma~n=?\]

OpenStudy (courtney_celiene1236):

Isn't that like the summation notation of n?

OpenStudy (sshayer):

\[\Sigma~n^2=\frac{ n \left( n+1 \right)\left( 2n+1 \right) }{ 6 }\] \[\Sigma n=\frac{ n \left( n+1 \right) }{ 2 }\]

OpenStudy (sshayer):

apply these and simplify if possible.

OpenStudy (sshayer):

can you show me your work?

OpenStudy (courtney_celiene1236):

I haven't done any work. I don't know anything about this, so I don't know what I'm doing.

OpenStudy (sshayer):

o kay i will solve.

OpenStudy (courtney_celiene1236):

I would like to understand how to do it tho.

OpenStudy (sshayer):

\[\sum t _{n}=16\frac{ n \left( n+1 \right)\left( 2n+1 \right) }{ 6 }+8\frac{ n \left( n+1 \right) }{ 2 }\]

OpenStudy (sshayer):

\[=8\frac{ n \left( n+1 \right) }{ 2 }\left[ \frac{ 2 }{ 3 } \left( 2n+1 \right)+1\right]\] \[=4n(n+1)\left[ \frac{ 4n+2+3 }{ 3 } \right]=\frac{ 1 }{ 3 }4n \left( n+1 \right)\left( 4n+5 \right)\]

OpenStudy (courtney_celiene1236):

I don't really understand.

OpenStudy (courtney_celiene1236):

@sleepyjess could you explain this?

OpenStudy (sshayer):

can you tell me where you have problem?

OpenStudy (courtney_celiene1236):

i don't understand how you got this

OpenStudy (sshayer):

(x+1)^2=x^2+2x+1 (x+1)^2-x^2=2x+1 put x=1,2,3,...,n 

OpenStudy (sshayer):



OpenStudy (courtney_celiene1236):

Thanks I think I kind of understand now.